# Using Biot-Savart to shrink Eulerian domains while maintaining or improving external flow accuracy

### v1

@@ Abstract
We introduce a novel boundary condition formulated using a Biot-Savart vorticity integral inside an Eulerian fluid domain that maintains high-accuracy results even when the domain boundary is within a body-length of immersed solid boundaries. The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step. We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded by the total vorticity contained in the lower levels and that it scales with the inverse power of the problem’s dimension. We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly. We show that the new method exactly captures the analytical added-mass force of a square plate accelerated from rest even when the entire domain only extends 1/2 diameter from the plate. The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 10 times larger to converge on the new method’s result. Finally, we study the sensitivity of deflected wakes to the new domain boundary conditions and show that only in minimal domains, 5 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body. Doubling the domain size recovers the correct results.
@@
External flow boundary condition via the Biot-Savart integral enables minimal computational domains
@@
Reduction in degrees of freedom by $O(10^{3})$ compared to reflection conditions for flow with bounded vorticity.
@@ 1 Introduction
Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain. Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain. While the former does not usually introduce significant errors, the latter can significantly impact the numerical solution when the finite domain is not sufficiently large. In the context of external flows, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence the forces experienced by the body [1].
Numerical simulations of external incompressible flows enforce far-field boundary conditions on the computational domain’s exterior. These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [2]. These can lead to significant errors. In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large. While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process.
Grid stretching away from the immersed body can efficiently maximize the domain size while maintaining a high-resolution mesh around the body, see [3, 4]. This can limit the blockage effects on the body. However, this method requires correct stretching ratios to avoid excessive numerical dissipation or reflection inside the domain. The method is also inefficient in cases where the body undergoes substantial motion within the computational domain, for example, when tracking the motion of a kite. Grid stretching can be replaced by a buffer layer that acts as a numerical damper which ensures that the solution decays rapidly before it reaches the boundaries [5]. Following a similar approach, a coarse grid can provide the boundary conditions for a relatively small, well-resolved domain; this is the basis of the multi-domain methods [1]. The method interpolates the coarse mesh solution to the walls of the finer domain, providing artificial far-field boundary conditions.
For flows with one [6, 7] or two unbounded directions [8], coordinate transforms or analytical potential flow solution can be applied in the unbounded directions and correctly capture solution. Potential flow boundary conditions are prescribed following the classical Hodge decomposition of the velocity field into an irrotational and an incompressible part
@@
where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively.
@@
Fundamental solution to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to approximate discrete differential equations on unbounded domains [9]. Points close to the source are evaluated with the discrete kernel (accurate) while points far away are evaluated with the asymptotic expansion (fast). These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [10] and flows with immersed-boundaries [11]. However, they are based on the solution of inhomogeneous, linear, constant coefficient elliptic equations and are thus incompatible with some immersed-boundary methods [3, 4].
Other methods leverage the strength of an Eulerian-based solver for the near body and Lagrangian solvers for the far wake by coupling a near wall Eulerian solver to a Vortex particle method in the wake [12]. They enable tracking the coarser wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface. However, several (user-defined) parameters are required to correctly interpolate the Vortex particle method solution onto the Eulerian mesh. Extension to three-dimensional flow is possible but will significantly increase the computational cost.
All these methods allow for a reduction of the error introduced by truncating an infinite computational domain. However, their generality is often limited to specific discretization of the governing equations or introduces complex coupling procedures.
@@
In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions on the velocity and pressure fields on the exterior of a Cartesian computational domain. The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the covering equations and is compatible with immersed boundaries. Additionally, it does not introduce additional Lagrangian or Eulerian mesh. In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations. Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field. We then demonstrate, with confined vorticity and vortex-street examples, that the new boundary conditions significantly outperform standard reflective boundary both in terms of computational time and solution accuracy. We present unexpected result that the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach. We close this manuscript with a discussion of the method’s applicability to various flows in science and engineering.
@@ 2 Method
We start by describing the governing equations and their numerical solution. The flow is governed by the Navier-Stokes equations for an incompressible flow. We discretize these equations on a staggered finite-volume mesh. The resulting system of equations is solved within a computational domain $\Omega(N)$, where $N$ is the number of degrees of freedom of the problem. The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem. In 2D we have $S=1/2$, and in 3D $S=2/3$. The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
@@
$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (2)
@@
Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. As discussed in the introduction, the simplest of these is a reflection condition
@@ 2.1 Biot-Savart boundary conditions
This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere. Such a formulation is exact for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [17]. However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
@@ 2.2 Fast evaluation of the Biot-Savart integral
The aim of this novel boundary conditions is to reduce the computational cost compared to a large grid by providing an accrate estimate of the far field conditions. A naive evaluation of Eq. 9  would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, overwhelming any savings from reducing the size of the domain. In vortex-particle methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations. Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
@@
where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2. In this work, we define nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$ and scale this width with the cell size which doubles at each level, i.e. $h^{(i)}=h^{(1)}2^{i-1}$. Therefore the domains are determined by $\vec{x}$ and a non-dimensional half-width
@@
Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate equation 12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
In practise, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation. The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
@@
as $\vec{e}_{z}$ is a unit vector. On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$. An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the $i$th-level evaluation of the Biot-Savart integral is therefore
@@ 2.3 Biot-Savart & Projection Algorithm
Substitution of the Biot-Savart boundary condition (Eq. 8) into the projection scheme results in a slightly modified intermediate velocity field
Biot-Savart intermediate step
$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (20)
$\vec{u}^{*}=f(\vec{x}_{b},\nabla\times\vec{u}^{*},\Omega)$ $\quad\forall\ \vec{x}_{b}\in\partial\Omega(N^{\cal S})$ (21)
where the exterior boundary condition have been computed from the intermediate velocity $\vec{u}^{*}$ in the domain. The projection step, however, results in a coupling between the body and the domain’s boundary through the pressure field
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}$ (22)
$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (23)
$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (24)
Where we have omitted the domain of definition of each equation for clarity. Away from the body, the source term in the Biot-Savart integral in Eq. 24  vanishes. However, due to the pressure boundary condition on the immersed body ($\partial p/\partial\hat{n}=\hat{n}\cdot(\nabla^{2}\vec{u}/Re-\text{D}\vec{U}_{b}/\text{D}t)$), a thin layer of vorticity is generated by the pressure gradient term that changes the induced velocity field on the outer domain’s boundary. The projection steps and the Biot-Savart boundary conditions now form a fixed-point system
Biot-Savart projection step
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}-\nabla\cdot f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (25)
$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (26)
$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (27)
This coupling only occurs on the domain’s boundary, and the resulting system is mildly non-linear. This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during each pressure solver iteration. This coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [19, 4].
Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed on the domain boundary. As such, we apply a uniform correction
$\vec{u}=\vec{u}-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (28)
ensuring a globally divergence-free field is passed to the pressure projection method.

### v2

# Using Biot-Savart boundary conditions for unbounded external flow on Eulerian meshes

@@ Abstract
We introduce a novel boundary condition formulated using a Biot-Savart vorticity integral inside an Eulerian fluid domain that maintains high-accuracy results even when the domain boundary is within a body-length of immersed solid boundaries. The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step. We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded and scales with the inverse power of the problem’s dimension. We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly. We show that the new method exactly captures the analytical added-mass force of accelerating plates and matches experimentally measured wake development even when the entire domain only extends 1/2 diameter from the plate. The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 48 times larger to converge on the new method’s result. Finally, we study the sensitivity of 2D deflected wakes to the new domain boundary conditions and show that only in minimal domains, 25 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body. Doubling the domain size recovers the correct results.
@@
Unbounded flow simulations on minimal Eulerian domains via Biot-Savart boundary conditions.
@@
Fully-couples velocity-pressure description to vorticity-condition via projection step.
Matches analytic and experimental results while reducing required domain size by 48x.
@@ 1 Introduction
Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain. Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain with far-field boundary conditions.
External incompressible flows enforce far-field boundary conditions on the computational domain’s exterior. These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [1]. These can lead to significant errors. In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large. Here, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence on the solution [2]. While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process. It is also computationally inefficient to add grid cells to fill such a large domain, and grid refinement near the solid boundaries introduces well-known sensitives to the grid expansion ratios and cell aspect ratios due to numerical diffusion, dispersion, reflection, and linear system stiffness. Alternatively, numerical dissipation can be explicitly imposed in a buffer layer on a smaller domain that forces the solution to decay rapidly before it reaches the boundaries [3].
Domain augmentation is a more principled approach to simulate external flow, such as using a large coarse grid to provide artificial far-field boundary conditions for a small well-resolved domain [2]. The outer flow field can also be augmented with a vortex particle-mesh method, reducing dispersive errors [4]. Such methods enable tracking the wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface and can also include immersed boundaries [5]. However, several (user-defined) parameters are required to correctly interpolate between the two domains, and the complexity of accurately coupling the domains makes theis apporach difficult to apply to existing flow solvers.
For flows with one [6] or two unbounded directions [7, 8], coordinate transforms and/or analytical potential flow solution [9] can be applied in the unbounded directions. Potential flow boundary conditions are prescribed following the classical Helmholtz decomposition of the velocity field into an irrotational and an incompressible part
@@
where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively. While these methods are elegant, they rely either on a single unbounded direction (as in [6] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [8]), rendering them difficult to generalize to 3D flow problems.
Other classes of methods solve unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [10] or by using the James-Lackner [11, 12] algorithm that solves two nested domain Poisson problems, the inner-most with homogeneous Neumann boundary conditions and second outermost domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator. To accelerate the convolution, multilevel approximation [13] or using local correction methods [14] can be used. Although these methods enforce unbounded boundary conditions, they require iterating a Poisson problem or solving it multiple times.
@@
Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations in unbounded domains [15, 16]. Points close to the source are evaluated with the discrete kernel (accurate), while points far away are evaluated with the asymptotic expansion (fast). These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [17] and flows with immersed-boundaries [18]. This allows for computational domains snug to the vorticity field, but adapting these methods to to existing Eulerian or immersed boundary solvers is difficult.
@@
In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions of a viscous flow simulation by coupling to a vorticity-based description of the flow on the boundary via the Biot-Savart equation. The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the governing equations, is compatible with immersed boundary methods and fast pressure projection methods, and does not require additional meshes to describe the flow. In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations. Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field. With confined vorticity and vortex wake examples, we demonstrate that the new boundary conditions significantly outperform standard reflective boundary conditions on significantly smaller domains. We present an unexpected result: the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and we use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach. We finish this manuscript by discussing the method’s applicability to various flows in science and engineering.
@@ 2 Method
We start by describing the governing equations and their numerical solution. The flow is governed by the Navier-Stokes equations for an incompressible flow. We discretize these equations on a staggered finite-volume mesh. The resulting system of equations is solved within a computational domain $\Omega(N)$, where $N$ is the number of degrees of freedom of the problem. The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem. In 2D we have ${\cal S}=1/2$, and in 3D ${\cal S}=2/3$. The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
@@
$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{\text{Re}}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (2)
@@
Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. The simplest of these is a reflection condition
@@ 2.1 Biot-Savart boundary conditions
This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere. This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [25]. However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
@@ 2.2 Fast evaluation of the Biot-Savart integral
The aim of this novel boundary conditions is to reduce the computational cost compared to a large grid by providing an accrate estimate of the far field conditions. A naive evaluation of Eq. 8 would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, potentially overwhelming any savings from reducing the size of the domain. In vortex particle-mesh methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations. Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
@@
where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2. In this work, we use nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$, and scale this width with the cell size which doubles at each level, i.e. $h^{(i)}=h^{(1)}2^{i-1}$. Therefore the domains are determined by $\vec{x}$ and a non-dimensional half-width
@@
Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate Eq. 12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
In practice, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation. The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
@@
as $\vec{e}_{z}$ is a unit vector. On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$. An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the multilevel evaluation of the Biot-Savart integral is therefore
@@ 2.3 Biot-Savart & Projection Algorithm
The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq. 2 and projecting it to be divergence free with Eq. 6. However, the new boundary condition introduces a coupling between the body and domain boundaries. To see this, we write the equation for the boundary velocity on $\vec{x}_{b}\in\partial\Omega(N^{\cal S})$ at the new time-step as
$\vec{u}^{t+\Delta t}(\vec{x}_{b})=f(\vec{x_{b}},\nabla\times\vec{u}^{t+\Delta t},\Omega)=f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)-f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)$ (20)
where we’ve substituted Eq. 6. The last term in Eq. 20 is non-zero on cells touching the body boundary $\partial\cal B$ because, while the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure boundary condition on the immersed body generates a thin layer of vorticity on $\partial\cal B$. This contribution is unknown until after the Poisson system of Eq. 5 is solved, but updating the normal velocity on $\partial\Omega(N^{\cal S})$ will change the source term for Eq. 5 on cells touching the domain boundary, requiring a new pressure solution.
This coupling between the domain and body boundaries results in a mildly non-linear fixed-point system. This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during pressure solver iteration. We use a Multi-Grid solver for the Poisson system, as described in [26], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle was sufficient to ensure the final pressure and velocity field satisfy both the momentum and divergence-free conditions without increasing the number of solver iterations. Note that this coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [27, 20]. This approach differs from the one in [10] for unbounded Poisson problems with embedded Dirichlet boundaries as we directly adjust the residual of the (iterative) Poisson solver and we do not iterate over the coupled Poisson/interior boundary problem. It is also different to the James-Lackner algorithm for the same reason [13].
Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed to machine precision on the domain boundary. As such, we apply a uniform correction to the velocity computed with Eq. 20
$\vec{u}(\vec{x}_{b})=\vec{u}(\vec{x}_{b})-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (21)
ensuring a globally divergence-free field is passed to the pressure projection method on each iteration.